OpenStudy (anonymous):

A baseball diamond is a square 90 feet on each side. Suppose the pitcher is in the exact middle of the diamond and the batter hits a ball to the first baseman. The pitcher needs to run to first base (to receive the throw from the first baseman) before the batter can get there. How much shorter is the pitcher's run to the base than the batter's?

OpenStudy (anonymous):



OpenStudy (anonymous):

@Gummibear127

OpenStudy (anonymous):

I know I need to use the Pythagorean theorem, just not how

OpenStudy (anonymous):

Oh lol i just learned that theorem give me a second

OpenStudy (anonymous):

Ok, this is an extra credit problem, so its harder (at least for me it is :P)

OpenStudy (anonymous):

*it's

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

There are not any options, its a show-your-work type of question

OpenStudy (anonymous):

45 ft.?

OpenStudy (anonymous):

That is what I was thinking too

OpenStudy (anonymous):

because it's about halfway inbetween each base

OpenStudy (anonymous):

hmm... don't I have to do a^2+b^2=c^2 and all that stuff?

OpenStudy (anonymous):

tha'ts what I was thinking

OpenStudy (anonymous):

but yes you do

OpenStudy (anonymous):

so you just divide 90 by 2

OpenStudy (anonymous):

pretty much

OpenStudy (anonymous):

so 90^2+90^2=90^2

OpenStudy (anonymous):

I'm almost positive that isn't the way to do it....

OpenStudy (anonymous):

ehh it was just an educated guess

OpenStudy (anonymous):

Ok, I will wait for someone who like if VERY confident and knows the answer exactly or can teach me it. Thanks though

OpenStudy (anonymous):

Im not really good with this maybe @iGreen can help

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

Can anyone help me out???

OpenStudy (igreen):

@amistre64

OpenStudy (anonymous):